How many grams of oxygen are in 3.01x10^23 particles of glucose


1
Expert's answer
2021-02-03T07:21:56-0500

Moles of glucose ( C6H12O6 )=

3.01 × 1023 / 6.02 × 1023 = 0.5 mol.

Number of moles of O = 0.5 × 6 = 3 mol .

Mass of O atom = 3 × 16 = 48 g .